# Boosting Text-to-Chart Retrieval with Semantic Insights

### v2

@@ 5.3. CRBench Overview
Overall, CRBench presents a more challenging and realistic benchmark for evaluating text-to-chart retrieval models.
5.4. Discussion
We discuss the quality and potential limitations of CRBench.
Crowdsourcing and Query Generation. While crowdsourcing was used to generate queries, we addressed quality issues through a second round of crowdsourcing and GPT-4o. This process helped ensure the relevance and clarity of the queries, resulting in a high-quality set of precise and fuzzy queries.
Limitations. Despite careful curation, crowdsourcing variability and the evolving landscape of data visualizations may require periodic updates to CRBench to maintain its relevance and accuracy.
6. Experiments
In this section, we conduct a set of experiments to evaluate whether the synthesized semantic insights can enhance text-to-chart retrieval performance across multiple benchmarks and settings. We also assess whether the ChartFinder model achieves state-of-the-art performance in text-to-chart retrieval tasks.
Table 2. Performance Comparison of Text-to-Chart Retrieval Models on CRBench Benchmark.
@@ 6. Experiments
VISTA (Zhou et al., 2024a) 196M 8.21 20.00 27.18 13.15 16.42 9.16 19.08 24.43 13.16 15.8
Univl-DR (Liu et al., 2022) 151M 4.10 12.82 15.38 7.84 9.65 3.82 10.69 14.50 6.90 8.70
MARVEL (Zhou et al., 2024b) 310M 1.54 5.13 7.69 3.08 4.63 0.00 0.76 1.53 0.46 2.23
CLIP (Radford et al., 2021) 151M 12.31 26.15 33.33 19.01 22.43 10.69 21.37 25.95 15.69 18.15
CLIP-DPR (Liu et al., 2022) 151M 9.74 18.46 23.59 13.60 15.95 5.34 13.74 16.79 8.51 10.48
EVA-CLIP (Sun et al., 2023) 149M 5.13 19.49 24.10 11.08 14.21 3.05 9.92 13.74 6.40 8.14
CoCa  (Cherti et al., 2023) 151M 8.72 21.54 25.1 14.11 16.76 3.05 8.40 9.92 5.19 6.33
Jina-CLIP-v2 (Koukounas et al., 2024) 865M 4.10 16.92 25.13 10.01 13.58 3.05 10.69 14.50 5.84 7.87
Long-CLIP-B (Zhang et al., 2025) 149M 26.67 53.85 62.56 37.99 43.90 22.90 40.46 51.91 30.29 35.33
@@
Long-CLIP-L (Zhang et al., 2025) 427M 41.03 75.90 79.49 55.32 55.32 38.93 67.18 74.81 50.99 56.77
@@
6.1. Experimental Settings
Datasets. We conduct experiments using three benchmarks: CRBench, VisText (Tang et al., 2023), and Chart-to-Text (Kantharaj et al., 2022). We compare the retrieval performance of different models on CRBench and perform zero-shot experiments on VisText and Chart-to-Text to evaluate the generalization of our model. The characteristics of these benchmarks are summarized in Table 1.
@@ 6.1. Experimental Settings
Metrics. To evaluate the effectiveness of the model in text-to-chart retrieval, we use the following metrics: Recall@1, Recall@5, Recall@10, MRR@10, and NDCG@10. These metrics are widely adopted in information retrieval tasks (Qin et al., 2020)
@@
Methods. We evaluate 10 methods in our experiments: (1) VISTA (Zhou et al., 2024a), (2) Univl-dr (Liu et al., 2022), (3) MARVEL (Zhou et al., 2024b), (4) CLIP (Radford et al., 2021), (5) CLIP-DPR (Liu et al., 2022), (6) EVA-CLIP (Sun et al., 2023), (7) CoCa (Cherti et al., 2023), (8) Jina-CLIP-v2 (Koukounas et al., 2024), (9) Long-CLIP (Zhang et al., 2025) (with Long-CLIP-B for the Base version and Long-CLIP-L for the Large version, which differ by the size of the vision encoder used). (10) Our Method: ChartFinder and ChartFinder-B: Both versions use the same Long-CLIP text encoder but differ in their visual encoder. ChartFinder uses the larger ViT-L/14, while ChartFinder-B uses the smaller ViT-B/16.
All methods encode the query and candidate charts into embeddings, with retrieval performed using cosine similarity via FAISS (Johnson et al., 2017).
Implementation Details. Many CLIP-based models apply center cropping during preprocessing to focus on the central part of the image. However, we observed that this operation often crops out crucial chart information, such as the title and axis labels ($x$-axis and $y$-axis names). This can negatively impact retrieval performance, introducing bias in the model’s ability to understand key chart features. To address this, we removed the center crop operation during testing and instead resized the images directly to a shape suitable for the model’s vision encoder input. This adjustment ensures that important chart elements remain intact, reducing potential bias and improving the retrieval process. We applied this image processing strategy consistently across all methods we evaluated.
@@
Table 3. Performance Comparison of Text-to-Chart Retrieval Models on VisText Benchmark.
@@
R@1 R@5 R@10 MRR@10 NDCG@10 R@1 R@5 R@10 MRR@10 NDCG@10
VISTA (Zhou et al., 2024a) 75.85 87.76 89.80 79.85 82.27 44.56 63.95 70.86 37.61 41.80
Univl-DR (Liu et al., 2022) 63.72 81.75 86.62 71.57 75.22 34.58 52.83 60.77 42.61 46.93
MARVEL (Zhou et al., 2024b) 47.51 66.55 72.79 55.61 59.73 24.04 39.80 48.64 31.25 35.37
CLIP  (Radford et al., 2021) 77.44 89.46 92.74 83.05 85.41 48.41 68.25 74.83 56.81 61.14
CLIP-DPR  (Liu et al., 2022) 74.26 88.10 90.93 80.59 83.14 45.01 64.74 72.11 53.51 57.97
EVA-CLIP (Sun et al., 2023) 71.77 87.07 89.68 78.50 81.26 31.86 49.43 57.26 39.39 43.64
CoCa  (Cherti et al., 2023) 75.62 84.35 87.87 79.69 81.65 44.22 58.96 66.44 50.54 54.30
Jina-CLIP-v2  (Koukounas et al., 2024) 79.71 90.93 94.67 84.85 87.23 54.20 74.49 80.50 62.72 67.01
Long-CLIP-B (Zhang et al., 2025) 90.48 94.78 95.35 81.88 83.50 60.20 78.12 83.45 67.99 71.73
ChartFinder-B 93.20 97.85 98.75 95.17 96.05 62.59 79.48 84.01 69.64 73.11
Long-CLIP-L (Zhang et al., 2025) 93.20 98.07 99.09 95.21 96.15 66.55 82.99 88.66 73.43 77.09
ChartFinder 96.15 99.43 99.66 96.90 97.44 67.80 84.58 89.12 75.20 78.78
Table 4. Experimental Results on Chart-to-Text Benchmark.
Models R@1 R@5 R@10 MRR@10 NDCG@10
VISTA  (Zhou et al., 2024a) 51.72 68.46 73.20 58.94 62.39
Univl-DR  (Liu et al., 2022) 49.61 67.91 73.08 57.60 61.35
MARVEL  (Zhou et al., 2024b) 21.68 38.19 46.30 28.97 33.08
CLIP  (Radford et al., 2021) 74.01 88.16 92.03 80.25 83.11
CLIP-DPR  (Liu et al., 2022) 66.04 82.56 85.86 73.09 76.21
EVA-CLIP (Sun et al., 2023) 68.05 83.42 87.80 74.73 77.90
CoCa  (Cherti et al., 2023) 67.98 83.92 87.65 74.49 77.92
Jina-CLIP-v2 (Koukounas et al., 2024) 65.83 83.78 88.23 73.51 77.08
Long-CLIP-B  (Zhang et al., 2025) 90.81 98.28 98.85 94.18 95.35
ChartFinder-B 94.90 99.07 99.43 96.89 97.53
Long-CLIP-L  (Zhang et al., 2025) 94.97 99.35 99.78 97.06 97.75
ChartFinder 97.06 99.86 99.86 98.41 98.80
6.2. Exp-1: Overall Results on CRBench
The purpose of this experiment is to evaluate and compare the performance of different models in precise and fuzzy query tasks.
Experimental results are shown in Table 2. From the experimental results, it can be seen that ChartFinder and ChartFinder-B perform best in both query tasks, and multiple metrics are ahead of other methods by more than 5% or even 10%. For example, in the precise query task, ChartFinder’s R@1, R@5, and R@10 are 47.18%, 80.00%, and 84.10%, respectively, significantly higher than the second-place Long-CLIP-L’s 41.03%, 75.90%, and 79.49%. In the fuzzy query task, ChartFinder also surpasses Long-CLIP-L, with R@1, R@5, and R@10 being 3.05%, 8.39%, and 5.34% higher, respectively. In comparison, although Jina-CLIP-v2 (865M) has the largest number of parameters, it performs poorly in both types of tasks. The R@1 in the precise query task is only 4.10%, and the R@1 in the fuzzy query task is also only 3.05%.
Overall, ChartFinder and ChartFinder-B perform significantly better in precise and fuzzy query tasks than all the other models, leading other methods by more than 5% or 10% in multiple metrics.
Finding 1: ChartFinder significantly outperforms existing methods on both precise and fuzzy queries. The performance gain is consistent across all the metrics.
6.3. Exp-2: ChartFinder in the Zero-Shot Setting
In this experiment, we evaluate the performance of ChartFinder in a zero-shot setting, where the model is tasked with retrieving relevant charts without any prior fine-tuning on the target dataset. This setup tests the model’s ability to generalize across different datasets, assessing how well ChartFinder can perform text-to-chart retrieval when exposed to new, unseen data.
6.3.1. Experiments on VisText Benchmark
We use the test set of VisText (Tang et al., 2023), comprising 882 charts with each chart corresponding to two captions mentioned in Section 5.3, to conduct this group of experiments. We take the two-level caption as the query, and the annotated chart as the retrieval target to perform text-to-chart retrieval tasks.
Overall Results. As shown in Table 3, ChartFinder performs best in all metricss in L1 caption, significantly ahead of other models. For example, Recall@1 reaches 96.15%, far exceeding other models. Recall@10 reaches 99.66%, close to a full match. MRR@10 and NDCG@10 are 96.90% and 97.44%, respectively, further verifying their advantage in ranking quality. ChartFinder-B still achieves excellent performance with fewer parameters. Recall@1 reaches 93.20%, which is comparable to Long-CLIP-L (93.20%) and about 3 percentage points higher than Long-CLIP-B. On Recall@10, ChartFinder-B reaches 98.85%, surpassing all models with similar parameter scales. L2+L3 captions are more inclined to detailed analytical descriptions. Thus, the retrieval difficulty is significantly higher than the L1 task, and the overall metrics have declined. ChartFinder still leads in this task with R@1 reaching 67.80%, R@10 reaching 89.12%, MRR@10 and NDCG@10 reaching 75.20% and 78.78%, respectively.
Finding 2: The results of all models decreased when L2+L3 was used as the query. We believe that the abstract data information, like a trend, cannot be fully understood by the model.
6.3.2. Experiments on Chart-to-Text Benchmark
We use the Chart-to-Text dataset (Kantharaj et al., 2022) to conduct this experiment. Specifically, this dataset has 1,393 text and chart pairs in the test set. We use the same metrics as the Section 6.3.1.
Overall Results. As shown in Table 4, ChartFinder significantly outperforms other models in all metrics. For example, Recall@1 reaches 97.06%, 2.09% higher than the second-best model (Long-CLIP-L, 94.97%). Recall@10 reaches 99.86%, almost close to the full score, showing the excellent performance of the model in the top-$k$ retrieval task. MRR@10 and NDCG@10 are 98.41% and 98.80%, respectively, further verifying the effectiveness of ChartFinder in ranking quality. ChartFinder-B also demonstrates obvious performance, surpassing Long-CLIP-B in all metrics, especially in Recall@1 (94.90% vs. 90.81%) and NDCG@10 (97.53% vs. 95.35%). The above results further prove the effectiveness of ChartFinder.
Finding 3: When we use long captions as the query, the accuracy of all models is improved compared to CRBench. This also shows that in real BI scenarios, retrieval is more difficult when the query is shorter and more brief..
6.4. Exp-3: Ablation Study on the Combination of Synthesized Semantic Insights
Experiment Settings. To verify the necessity and synergy of the three semantic insights we proposed, we designed a series of ablation experiments. We used ChartFinder-B as the base model and arranged and combined the three semantic insights to obtain a total of 7 training configurations (including the initial model). Table  5 shows the performance of each configuration in the two scenarios of precise query and fuzzy query, and the evaluation indicators include R@10, MRR@10, and NDCG@10.
Overall Results. The experimental results reveal several key findings: First, the introduction of any semantic insights can significantly improve the performance of the baseline model, among which statistics-oriented insights perform best with an overall performance of 48.35%, slightly better than task-oriented (48.20) and visual-oriented (46.67) insights.
In the dual combination, the combination of visual-oriented and task-oriented insights has the most outstanding effect (50.81). It is worth noting that although statistics-oriented insights perform best when used alone, they have limited benefits when combined with other types, which clearly shows that there are obstacles to information integration between different types of insights: the abstract numerical analysis provided by statistical insights is difficult to directly establish an effective connection with visual patterns or application scenarios.
However, when we use the three insights together, we get the best result (54.34). We believe that this is because the three types of insights form a complete information processing chain: from visual perception (visual-oriented) to statistical verification (statistics-oriented) to application transformation (task-oriented). When one link is missing, it is difficult for the model to establish an effective semantic connection, resulting in reduced information utilization efficiency.
Finding 4: The best result (54.34%) was achieved by combining all three multiple semantic insights because they form a complete information processing chain from visual perception to statistical verification to application conversion. Using one insight alone or combining only two insights has a limited effect, indicating that there are information integration barriers between different types of insights. This proves the effectiveness of the three semantic insights we proposed.
@@ 6.4. Exp-3: Ablation Study on the Combination of Synthesized Semantic Insights
Table 5. The Effectiveness of ChartFinder on CRBench by Incorporating Different Semantic Insights.
Models Visual-o Statistics-o Task-o Precise Query Fuzzy Query Overall
@@
6.5. Exp-4: The Generalizability of Synthesized Semantic Insights

### v4

@@ 5 Experiments
Table 2: Performance Comparison of Text-to-Chart Retrieval Models on CRBench Benchmark.
@@
VISTA Zhou et al. (2024a) 196M 8.21 20.00 27.18 13.15 16.42 9.16 19.08 24.43 13.16 15.8
Univl-DR Liu et al. (2022) 151M 4.10 12.82 15.38 7.84 9.65 3.82 10.69 14.50 6.90 8.70
MARVEL Zhou et al. (2024b) 310M 1.54 5.13 7.69 3.08 4.63 0.00 0.76 1.53 0.46 2.23
CLIP Radford et al. (2021) 151M 12.31 26.15 33.33 19.01 22.43 10.69 21.37 25.95 15.69 18.15
CLIP-DPR Liu et al. (2022) 151M 9.74 18.46 23.59 13.60 15.95 5.34 13.74 16.79 8.51 10.48
EVA-CLIP Sun et al. (2023) 149M 5.13 19.49 24.10 11.08 14.21 3.05 9.92 13.74 6.40 8.14
CoCa Cherti et al. (2023) 151M 8.72 21.54 25.1 14.11 16.76 3.05 8.40 9.92 5.19 6.33
Jina-CLIP-v2 Koukounas et al. (2024) 865M 4.10 16.92 25.13 10.01 13.58 3.05 10.69 14.50 5.84 7.87
Long-CLIP-B Zhang et al. (2025) 149M 26.67 53.85 62.56 37.99 43.90 22.90 40.46 51.91 30.29 35.33
@@
Long-CLIP-L Zhang et al. (2025) 427M 41.03 75.90 79.49 55.32 55.32 38.93 67.18 74.81 50.99 56.77
@@
Metrics. To evaluate the effectiveness of the model in text-to-chart retrieval, we use the following metrics: Recall@1, Recall@5, Recall@10, MRR@10, and NDCG@10. These metrics are widely adopted in information retrieval tasks Qin et al. (2020)
@@
Methods. We evaluate 10 methods in our experiments: (1) VISTA Zhou et al. (2024a), (2) Univl-dr Liu et al. (2022), (3) MARVEL Zhou et al. (2024b), (4) CLIP Radford et al. (2021), (5) CLIP-DPR Liu et al. (2022), (6) EVA-CLIP Sun et al. (2023), (7) CoCa Cherti et al. (2023), (8) Jina-CLIP-v2 Koukounas et al. (2024), (9) Long-CLIP Zhang et al. (2025) (with Long-CLIP-B for the Base version and Long-CLIP-L for the Large version, which differ by the size of the vision encoder used). (10) Our Method: ChartFinder and ChartFinder-B: Both versions use the same Long-CLIP text encoder but differ in their visual encoder. ChartFinder uses the larger ViT-L/14, while ChartFinder-B uses the smaller ViT-B/16. In all experiments, we changed the image preprocessing operation from center crop to direct resize, because we found that cropping would crop important information such as the title of the image.
Training Details. We trained ChartFinder using 4 A800 GPUs with a learning rate of 1e61e-6 for 20 epochs. For the base version of the model, we set the batch size to 512, while for the larger version, we reduced the batch size to 256 due to the increased memory usage associated with the larger model.
5.1 Exp-1: Overall Results on CRBench
Experiment Settings. The purpose of this experiment is to evaluate and compare the performance of different models in precise and fuzzy query tasks.
Overall Results.Experimental results are shown in Table 2. We can observe that ChartFinder and ChartFinder-B perform best in both query tasks, and multiple metrics are ahead of other methods by more than 5% or even 10%. For example, in the precise query task, ChartFinder’s R@1, R@5, and R@10 are 47.18%, 80.00%, and 84.10%, respectively, significantly higher than the second-place Long-CLIP-L’s 41.03%, 75.90%, and 79.49%. In the fuzzy query task, ChartFinder also surpasses Long-CLIP-L, with R@1, R@5, and R@10 being 3.05%, 8.39%, and 5.34% higher, respectively. In comparison, although Jina-CLIP-v2 (865M) has the largest number of parameters, it performs poorly in both types of tasks. The R@1 in the precise query task is only 4.10%, and the R@1 in the fuzzy query task is also only 3.05%.
6 Limitations
Dependence and Scope of Semantic Insight Generation. The insight generation process utilizes a LLM(LLaMA-3.1), meaning the quality of the output is subject to the model’s inherent knowledge and potential biases. Although the study validated a small sample of insights via crowdsourcing, it cannot guarantee that all 207,498 synthesized insights are entirely free of errors or misinterpretations.
Limitations of the CRBench Benchmark Construction. The queries in CRBench were generated by GPT-4o, not from organic user query logs. While this approach was designed to simulate realistic precise and fuzzy queries and was validated through a rigorous crowdsourcing process, these synthetic queries may not fully capture the diversity, ambiguity, and unforeseen intents of genuine user behavior.
Model Architecture and Application Scenario Constraints. The core architecture of ChartFinder is based on CLIP. Although the training strategy is innovative, its fundamental performance is still bound by the inherent capabilities and limitations of the underlying vision-language models.
Generalizability of Data Sources . The datasets used for generating the training data were sourced exclusively from Kaggle. While these are high-quality and derived from real-world applications, they may not fully represent the entire spectrum of data types and structures found within private, enterprise Business Intelligence (BI) systems, which can be more domain-specific or messy.
Appendix A Complete Experiments
A.1 Exp-2: ChartFinder in the Zero-Shot Setting
To evaluate the generalization ability of ChartFinder, we conducted zero-shot experiments on two external public benchmarks, VisText Tang et al. (2023) and Chart-to-Text Kantharaj et al. (2022). In this setting, we directly applied the model trained on our synthetic data to these new tasks without any fine-tuning to test the transferability of its learned semantic understanding ability.
A.1.1 Experiments on VisText Benchmark.
We use the test set of VisText Tang et al. (2023), comprising 882 charts with each chart corresponding to two captions mentioned in Section 3.3, to conduct this group of experiments. We take the two-level caption as the query, and the annotated chart as the retrieval target to perform text-to-chart retrieval tasks.
Overall Results. As shown in Table 3, ChartFinder performs best in all metricss in L1 caption, significantly ahead of other models. For example, Recall@1 reaches 96.15%, far exceeding other models. Recall@10 reaches 99.66%, close to a full match. MRR@10 and NDCG@10 are 96.90% and 97.44%, respectively, further verifying their advantage in ranking quality. ChartFinder-B still achieves excellent performance with fewer parameters. Recall@1 reaches 93.20%, which is comparable to Long-CLIP-L and about 3 percentage points higher than Long-CLIP-B. On Recall@10, ChartFinder-B reaches 98.85%, surpassing all models with similar parameter scales. L2+L3 captions are more inclined to detailed analytical descriptions. Thus, the retrieval difficulty is significantly higher than the L1 task, and the overall metrics have declined. ChartFinder still leads in this task with R@1 reaching 67.80%, R@10 reaching 89.12%, MRR@10 and NDCG@10 reaching 75.20% and 78.78%, respectively.
A.1.2 Experiments on Chart-to-Text Benchmark
We use the Chart-to-Text dataset Kantharaj et al. (2022) to conduct this experiment. Specifically, this dataset has 1,393 text and chart pairs in the test set. We use the same metrics as other experiments.
Overall Results. As shown in Table 5, ChartFinder outperforms other models in all metrics. For example, Recall@1 reaches 97.06%, 2.09% higher than the second-best model (Long-CLIP-L, 94.97%). Recall@10 reaches 99.86%, almost close to the full score, showing the excellent performance of the model in the top-$k$ retrieval task. MRR@10 and NDCG@10 are 98.41% and 98.80%, respectively, further verifying the effectiveness of Liter in ranking quality. ChartFinder-B also demonstrates obvious performance, surpassing Long-CLIP-B in all metrics, especially in Recall@1 (94.90% vs. 90.81%) and NDCG@10 (97.53% vs. 95.35%).
Finding 1: When we use long captions as the query, the accuracy of all models is improved compared to CRBench. This also shows that in real BI scenarios, retrieval is more difficult when the query is shorter and more brief.
Although the Chart-to-Text dataset can effectively test the performance of the model in clear semantic scenarios, its task design also has certain limitations. For example, the same pictures exist in the dataset, which narrows the search scope. Most of the queries in Chart-to-Text are descriptive sentences generated based on charts. There is a lack of user intent-driven queries, which is not enough to fully evaluate the generalization and upper-limit capabilities of the model. In contrast, the queries in CRBench are more complex and closer to actual applications, which can better test the performance of the model in real retrieval scenarios.
@@ A.1.2 Experiments on Chart-to-Text Benchmark
Table 3: Performance Comparison of Text-to-Chart Retrieval Models on VisText Benchmark.
@@
R@5 R@10 MRR@10 NDCG@10 R@5 R@10 MRR@10 NDCG@10
VISTA 87.76 89.80 79.85 82.27 63.95 70.86 37.61 41.80
Univl-DR 81.75 86.62 71.57 75.22 52.83 60.77 42.61 46.93
MARVEL 66.55 72.79 55.61 59.73 39.80 48.64 31.25 35.37
CLIP 89.46 92.74 83.05 85.41 68.25 74.83 56.81 61.14
CLIP-DPR 88.10 90.93 80.59 83.14 64.74 72.11 53.51 57.97
EVA-CLIP 87.07 89.68 78.50 81.26 49.43 57.26 39.39 43.64
CoCa 84.35 87.87 79.69 81.65 58.96 66.44 50.54 54.30
Jina-CLIP-v2 90.93 94.67 84.85 87.23 74.49 80.50 62.72 67.01
Long-CLIP-B 94.78 95.35 81.88 83.50 78.12 83.45 67.99 71.73
ChartFinder-B 97.85 98.75 95.17 96.05 79.48 84.01 69.64 73.11
Long-CLIP-L 98.07 99.09 95.21 96.15 82.99 88.66 73.43 77.09
ChartFinder 99.43 99.66 96.90 97.44 84.58 89.12 75.20 78.78
@@
Table 4: The Effectiveness of ChartFinder on CRBench by Incorporating Different Semantic Insights.
Visual-o Statistics-o Task-o Precise Query Fuzzy Query Overall
@@
Table 5: Experimental Results on Chart-to-Text Benchmark.
Models R@5 R@10 MRR@10 NDCG@10
VISTA 68.46 73.20 58.94 62.39
Univl-DR 67.91 73.08 57.60 61.35
MARVEL 38.19 46.30 28.97 33.08
CLIP 88.16 92.03 80.25 83.11
CLIP-DPR 82.56 85.86 73.09 76.21
EVA-CLIP 83.42 87.80 74.73 77.90
CoCa 83.92 87.65 74.49 77.92
Jina-CLIP-v2 83.78 88.23 73.51 77.08
Long-CLIP-B 98.28 98.85 94.18 95.35
ChartFinder-B 99.07 99.43 96.89 97.53
Long-CLIP-L 99.35 99.78 97.06 97.75
ChartFinder 99.86 99.86 98.41 98.80
A.2 Exp-3: Ablation Study on the Combination of Synthesized Semantic Insights
Experiment Settings. To verify the necessity and synergy of the three semantic insights we proposed, we designed a series of ablation experiments. We used ChartFinder-B as the base model and arranged and combined the three semantic insights to obtain a total of 7 training configurations (including the initial model). Table 4 shows the performance of each configuration in the two scenarios of precise query and fuzzy query, and the evaluation indicators include R@10, MRR@10, and NDCG@10.
Overall Results. The ablation study in Table 4 reveals several key insights. First, adding any single type of semantic insight significantly improves performance over the baseline, with statistics-oriented insights providing the largest individual boost (from 43.66 to 48.35). Second, combining two types of insights generally yields further improvement, with the visual-task combination being the most effective pair (50.81). However, the greatest performance leap occurs when all three insights are used together, achieving the top overall score of 54.34.
The experiments demonstrate that visual-oriented insights primarily assist in capturing general visual patterns, while statistics-oriented insights enhance the model’s analytical sensitivity to underlying data trends. Task-oriented insights further contextualize these representations, significantly improving retrieval precision in complex scenarios. Therefore, the optimal performance achieved by combining all three insights highlights their complementary nature and synergistic effects.
This suggests that while each insight is valuable, they form a complete information processing chain from visual perception (visual-oriented) to data-driven analysis (statistics-oriented) and finally to practical application (task-oriented). The full combination provides a holistic semantic context that allows the model to learn the most effective representations, demonstrating a powerful synergistic effect that surpasses any partial combination.
Finding 2: The three semantic insights are not merely additive but synergistic, with the complete set being crucial for optimal performance.
Table 6: Experiment results on Text-to-chart retrieval task.
Models Precise Query Fuzzy Query
R@1 R@5 R@10 MRR@10 NDCG@10 R@1 R@5 R@10 MRR@10 NDCG@10
CLIP-DPR 4.10 15.38 20.00 8.90 11.55 5.34 7.63 10.69 6.60 7.54
FT-CLIP-DPR 8.21 19.49 24.62 13.08 15.80 7.63 14.50 19.08 10.97 12.88
$\uparrow$4.11 $\uparrow$4.11 $\uparrow$4.62 $\uparrow$4.18 $\uparrow$4.25 $\uparrow$2.29 $\uparrow$6.87 $\uparrow$8.39 $\uparrow$4.37 $\uparrow$5.34
UniVL-DR 3.08 11.79 17.44 6.59 9.13 5.34 6.87 9.92 6.26 7.11
FT-UniVL-DR 8.21 17.95 26.15 13.25 16.29 12.21 18.32 20.61 15.08 16.42
$\uparrow$5.13 $\uparrow$6.16 $\uparrow$8.71 $\uparrow$6.66 $\uparrow$7.16 $\uparrow$6.87 $\uparrow$11.45 $\uparrow$10.69 $\uparrow$8.82 $\uparrow$9.31
Long-CLIP-B 26.67 53.85 62.56 37.99 43.90 22.90 40.46 51.91 30.29 35.33
ChartFinder-B 38.97 62.05 70.26 49.30 61.30 29.77 50.38 61.83 38.96 44.41
$\uparrow$12.30 $\uparrow$8.20 $\uparrow$7.70 $\uparrow$11.31 $\uparrow$17.40 $\uparrow$6.87 $\uparrow$9.92 $\uparrow$9.92 $\uparrow$8.67 $\uparrow$9.08
Long-CLIP-L 41.03 75.90 79.49 55.32 55.32 38.93 67.18 74.81 50.99 56.77
ChartFinder 47.18 80.00 84.10 61.23 66.90 41.98 75.57 80.15 55.31 61.40
$\uparrow$6.15 $\uparrow$4.10 $\uparrow$4.61 $\uparrow$5.91 $\uparrow$11.58 $\uparrow$3.05 $\uparrow$8.39 $\uparrow$5.34 $\uparrow$4.32 $\uparrow$4.63
A.3 Exp-4: The Generalizability of Synthesized Semantic Insights